OpenStudy (anonymous):

The only forces acting on a 2.7 kg mass are ~ F1=F1x ˆi−F1y ^j and F2=F2x ˆi+F2y ˆj, where F1x = 2 N, F1y = 7 N, F2x = 6 N, F2y = 2 N. What is the magnitude a of the acceleration of the particle? Answer in units of m/s2

OpenStudy (ghazi):

can you find the resultant force?

OpenStudy (anonymous):

by doing cos and sin for each?

OpenStudy (ghazi):

no, by using vectors

OpenStudy (anonymous):

Do i solve for the x and y component for each?

OpenStudy (ghazi):

\[F _{1}=(2 i-7j ) N\] and \[F _{2}=(6i+2j) N\] now resultant vector is \[F=F _{1}+F _{2}=(2i-7j)+(6i+2j)=(8i-5j) N\] and magnitude is of \[F _{R}=\sqrt{8^2+5^2}=\sqrt{64+25}=\sqrt{89}=9.43N\] therefore acceleration = 9.43/2.7 m/s^2

OpenStudy (ghazi):

@ncarrillo52 does that help?

OpenStudy (anonymous):

yes thank you!!

OpenStudy (ghazi):

:)

OpenStudy (anonymous):

Consider the three blocks connected by massless inextensible cords over massless frictionless pulleys as shown in the figure below. The coefficient of static friction between the (middle) block and the table top is µs = 0.54 and the system is in equilibrium. The acceleration of gravity is 9.8 m/s  SOLVE THE MAX WEIGHT M COULD BE.

OpenStudy (anonymous):

NVM I figured that one out. i forgot to time friction by normal.

OpenStudy (ghazi):

friction force= u*mass*acceleration due to gravity